# ATLAS: Automated Toolkit for Large-Scale Verified Code Synthesis

### v1

@@ 3.2.2. DafnySynthesis
GPT-4111As reported in (DafnySynthesis) at the time of benchmark release. 33.1 49.4 53.4
@@
Figure 7. Evaluation results on DafnySynthesis.
To evaluate our model on Dafny code synthesis, we use the Dafny Synthesis benchmark (DafnySynthesis), which consists of 228 programming tasks adapted for Dafny verification split into a test set of 178 tasks and 50 tasks for training. This dataset derives from the sanitized MBPP (Mostly Basic Python Programming) dataset (mbpp), curated to exclude problems requiring Python-specific features or external libraries incompatible with Dafny’s standard libraries. Each problem includes a natural language description of the task, Dafny method signature, and 3 test cases. The evaluation of the benchmark considers a response successful if it verifies without errors and contains non-trivial specifications of the functions or methods synthesized.
We evaluate our model using the signature prompt format from (DafnySynthesis), which combines the problem description with its method signature and test cases. While (DafnySynthesis) explored contextless and Chain of Thought (CoT) prompting, we selected the signature format as it provides a standardized and practical assessment of verification-aware programming, balancing realistic context with minimal scaffolding.
Results are shown in Figure 7. ATLAS-trained Qwen 2.5 Coder 7B achieves 65.8% verification success at Pass@5, surpassing GPT-4 (53.4%) by 12.4 percentage points, which was the best performing model at benchmark release. The substantial improvement over the baseline model Qwen 2.5 Coder 7B (15.8% at Pass@5) demonstrates the effectiveness of the ATLAS dataset in inducing verification-aware programming capabilities.
@@
4. Related and Future Work
Complementary to our work, dafny-annotator (dafny-annotator) targets automated annotation generation by bootstrapping from existing Dafny programs. In contrast, ATLAS employs a stage-wise approach that generates specifications before implementations without requiring Dafny code as input. While dafny-annotator evaluates on a test split of DafnyBench after using the remainder for synthesis and training, ATLAS achieves strong performance on DafnyBench without in-distribution training, demonstrating that general-purpose synthesis can effectively support specialized verification tasks.
Several promising research directions remain unexplored. First, while current benchmarks evaluate free-form synthesis or specialized tasks like proof hint infilling, there is a need for benchmarks emphasizing end-to-end synthesis for complete specifications. Recent work such as (vericoding), VERINA (verina), and CLEVER (clever) in Lean represent progress in this direction.
Second, while reinforcement learning has shown success in mathematical theorem proving (wang2025kiminaproverpreviewlargeformal; lin2025goedelproverfrontiermodelopensource), its application to verification-aware code generation remains underexplored due to lack of well-defined reward structures. Future work should investigate combining compiler feedback, complexity analysis, and proof resource allocation as reward signals.
Finally, recent advances in agentic theorem proving, including HILBERT (hilbert) and SeedProver (seed-prover), suggest promising directions. These approaches demonstrate how agentic interaction with verification frameworks can augment model capabilities through iterative refinement and guided exploration, with potential for practical co-pilot and agent tools.
@@ Appendix A Appendix
1:procedure GenerateLemmas($m,\sigma,R,E,\mathcal{T}$)
2:  Input: Method $m$, signature $\sigma$, pre-conditions $R$,
3:    post-conditions $E$, test cases $\mathcal{T}=\{(a_{i},o_{i},\Gamma_{i})\}_{i=1}^{n}$
4:    ($a_{i}$: inputs, $o_{i}$: outputs, $\Gamma_{i}$: input variable bindings)
5:  Output: Set of verification lemmas $\mathcal{L}$
6:  $\mathcal{L}\leftarrow\emptyset$
7:  for $i=1$ to $n$ do
8:   $(a_{i},o_{i},\Gamma_{i})\leftarrow\mathcal{T}_{i}$
9:   // Soundness lemma
10:   $R_{i}\leftarrow R\cup\Gamma_{i}\cup\{\sigma=a_{i}\}\cup\{v_{\text{out}}=o_{i}\}$
11:   $\ell^{\text{sound}}_{i}\leftarrow\text{ConstructLemma}(\sigma,v_{\text{out}},R_{i},E)$
12:   $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell^{\text{sound}}_{i}\}$
13:   // Completeness lemma (contradiction)
14:   $R_{i}\leftarrow R\cup\Gamma_{i}\cup\{\sigma=a_{i}\}\cup\{v_{\text{out}}\neq o_{i}\}\cup E$
15:   $\ell^{\text{contra}}_{i}\leftarrow\text{ConstructLemma}(\sigma,v_{\text{out}},R_{i},\{\bot\})$
16:   $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell^{\text{contra}}_{i}\}$
17:   // Completeness lemma (perturbation)
18:   $p_{i}\leftarrow\text{Perturbation}(o_{i},E)$
19:   $R_{i}\leftarrow R\cup\Gamma_{i}\cup\{\sigma=a_{i}\}\cup\{v_{\text{out}}=p_{i}\}$
20:   $\ell^{\text{perturb}}_{i}\leftarrow\text{ConstructLemma}(\sigma,v_{\text{out}},R_{i},E)$
21:   $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell^{\text{perturb}}_{i}\}$
22:  end for
23:  return $\mathcal{L}$
24:end procedure

### v2

@@ 4.2.2 DafnySynthesis
GPT-4111As reported in (Misu et al., 2024b) at the time of benchmark release. 33.1 49.4 53.4
@@
Figure 9: Evaluation results on DafnySynthesis.
To evaluate our model on Dafny code synthesis, we use the DafnySynthesis benchmark (Misu et al., 2024b), which consists of 228 programming tasks adapted for Dafny verification split into a test set of 178 tasks and 50 tasks for training. This dataset derives from the sanitized MBPP (Mostly Basic Python Programming) dataset (Austin et al., 2021), curated to exclude problems requiring Python-specific features or external libraries incompatible with Dafny’s standard libraries. Each problem includes a natural language description of the task, Dafny method signature, and 3 test cases. The evaluation of the benchmark considers a response successful if it verifies without errors and contains non-trivial specifications of the functions or methods synthesized.
We evaluate our model using the signature prompt format from (Misu et al., 2024b), which combines the problem description with its method signature and test cases. While (Misu et al., 2024b) explored contextless and Chain of Thought (CoT) prompting, we selected the signature format as it provides a standardized and practical assessment of verification-aware programming, balancing realistic context with minimal scaffolding.
Results are shown in Figure˜9. ATLAS-trained Qwen 2.5 Coder 7B achieves 65.8% verification success at Pass@5, surpassing GPT-4 (53.4%) by 12.4 percentage points, which was the best performing model at benchmark release. The substantial improvement over the baseline model Qwen 2.5 Coder 7B (15.8% at Pass@5) demonstrates the effectiveness of the ATLAS dataset in inducing verification-aware programming capabilities.
@@
5 Related and Future Work
Complementary to our work, dafny-annotator (Poesia et al., 2024) targets automated annotation generation by bootstrapping from existing Dafny programs. In contrast, ATLAS employs a stage-wise approach that generates specifications before implementations without requiring Dafny code as input. While dafny-annotator evaluates on a test split of DafnyBench after using the remainder for synthesis and training, ATLAS achieves strong performance on DafnyBench without in-distribution training, demonstrating that general-purpose synthesis can effectively support specialized verification tasks.
Several promising research directions remain unexplored. First, while current benchmarks evaluate free-form synthesis or specialized tasks like proof hint infilling, there is a need for benchmarks emphasizing end-to-end synthesis for complete specifications. Recent work such as (Bursuc et al., 2025), VERINA (Ye et al., 2025), and CLEVER (Thakur et al., 2025) in Lean represent progress in this direction.
Second, while reinforcement learning has shown success in mathematical theorem proving (Wang et al., 2025; Lin et al., 2025), its application to verification-aware code generation remains underexplored due to lack of well-defined reward structures. Future work should investigate combining compiler feedback, complexity analysis, and proof resource allocation as reward signals.
Finally, recent advances in agentic theorem proving, including HILBERT (Varambally et al., 2025) and SeedProver (Chen et al., 2025), suggest promising directions. These approaches demonstrate how agentic interaction with verification frameworks can augment model capabilities through iterative refinement and guided exploration, with potential for practical co-pilot and agent tools.
6 Conclusion
Our work ATLAS, an automated pipeline for synthesizing verified Dafny programs from algorithmic problems with reference implementations. By introducing soundness and completeness checks that leverage test cases, ATLAS ensures specification quality and filters degenerate solutions. Applied to TACO-verified, the pipeline produces 2.7K verified programs, which we decompose into 19K training examples spanning synthesis, repair, and proof infilling tasks. Fine-tuning Qwen 2.5 7B Coder on this data yields substantial improvements on both DafnyBench (32.4% to 56.9%) and DafnySynthesis (15.8% to 65.8%), demonstrating that synthetic data generation is a viable path to scaling LLM capabilities for formal verification. Our results suggest that the data scarcity bottleneck for verification-aware languages can be addressed through automated synthesis pipelines, opening opportunities for broader adoption of formal methods in software development.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
Soundness and Completeness Lemma Generation
Algorithm 1 presents the procedure for constructing verification lemmas that evaluate specification quality. The algorithm takes as input a method $m$ with its signature $\sigma$, pre-conditions $R$, post-conditions $E$, and a test suite $\mathcal{T}$ containing $n$ test cases. Each test case $(a_{i},o_{i},\Gamma_{i})$ comprises input arguments $a_{i}$, expected output $o_{i}$, and input variable bindings $\Gamma_{i}$ that map method parameters to concrete values.
Soundness Lemmas.
We instantiate the contract with concrete test inputs and outputs as preconditions, then verify that the postconditions hold. Verification failure indicates the specification is inconsistent with known correct behavior.
Contradiction Lemmas.
We assume the contract holds with a negated output and attempt to derive false. Success proves the postconditions uniquely determine the output, though such proofs often exceed automated prover capabilities.
Perturbation Lemmas.
We substitute LLM-generated incorrect outputs—such as negated booleans, offset integers, or sequences with swapped elements—and check if the contract still holds. Successful verification reveals the specification admits incorrect behavior and requires strengthening.
Feedback Integration.
The complete lemma set $\mathcal{L}$ is submitted to the Dafny verifier. Verification results are parsed and formatted into structured feedback:
Soundness Check Results:
Test 1: PASS (contract consistent with test case)
Test 2: FAIL (contract rejects valid input-output pair)
Completeness Check Results:
Test 1: PASS (perturbed output correctly rejected)
Test 2: FAIL (perturbed output incorrectly accepted)
This feedback is provided to the LLM Judge, which synthesizes actionable recommendations for specification refinement. Common remediation patterns include strengthening postconditions with additional constraints, introducing auxiliary predicates to capture invariant properties, and adding quantified assertions over collection elements.
@@
1:Input: Method $m$, signature $\sigma$, pre-conditions $R$, post-conditions $E$, test cases $\mathcal{T}=\{(a_{i},o_{i},\Gamma_{i})\}_{i=1}^{n}$ ($a_{i}$: inputs, $o_{i}$: outputs, $\Gamma_{i}$: input variable bindings)
2:Output: Set of verification lemmas $\mathcal{L}$
[…]
4:for $i=1$ to $n$ do
5:  $(a_{i},o_{i},\Gamma_{i})\leftarrow\mathcal{T}_{i}$
6:  // Soundness lemma
7:  $R_{i}\leftarrow R\cup\Gamma_{i}\cup\{\sigma=a_{i}\}\cup\{v_{\text{out}}=o_{i}\}$
8:  $\ell^{\text{sound}}_{i}\leftarrow\text{ConstructLemma}(\sigma,v_{\text{out}},R_{i},E)$
9:  $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell^{\text{sound}}_{i}\}$
10:  // Completeness lemma (contradiction)
11:  $R_{i}\leftarrow R\cup\Gamma_{i}\cup\{\sigma=a_{i}\}\cup\{v_{\text{out}}\neq o_{i}\}\cup E$
12:  $\ell^{\text{contra}}_{i}\leftarrow\text{ConstructLemma}(\sigma,v_{\text{out}},R_{i},\{\bot\})$
13:  $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell^{\text{contra}}_{i}\}$
14:  // Completeness lemma (perturbation)
15:  $p_{i}\leftarrow\text{Perturbation}(o_{i},E)$
16:  $R_{i}\leftarrow R\cup\Gamma_{i}\cup\{\sigma=a_{i}\}\cup\{v_{\text{out}}=p_{i}\}$
17:  $\ell^{\text{perturb}}_{i}\leftarrow\text{ConstructLemma}(\sigma,v_{\text{out}},R_{i},E)$
18:  $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell^{\text{perturb}}_{i}\}$
19:end for
20:return $\mathcal{L}$